A brick lies perilously close to the edge of the flat roof of a building. The roof edge is 50 ft above street level, and the brick has 270.0 J of potential energy with respect to street level.
a. Someone edges the brick off the roof, and it begins to fall. What is the brick’s kinetic energy when it is 35 ft above street level?
b. What is the brick’s kinetic energy the instant before it hits the street surface?
1
Expert's answer
2019-12-04T07:09:44-0500

"50ft.=15.24 m"

"35ft.=10.67m"

Potential energy

"=mgh=m*9.81*15.24=270J"

"\implies m= 1.8kg"


a) Initial kinetic energy is 0.

Using conservation of energy;

"270+0=K_f+mgh_f"

"K_f=270-(1.8*9.81*10.67)"

"=81.59J" (Answer)


b) The instant before the brick hits the ground, it's potential energy becomes 0.

Thus, all of its initial energy is converted to kinetic energy.

Hence, final kinetic energy"=270J." (Answer)